# Mitigating Covariate Shift in Imitation Learning for Autonomous Vehicles Using Latent Space Generative World Models

### v2

@@ IV-D DRIVE Sim Experiments
We also tested our model in NVIDIA DRIVE Sim, which is more photo-realistic and physically realistic than CARLA (Fig. 5). We trained our system using a 450-hour internal real driving dataset collected by multiple vehicles in our autonomous vehicle fleet in several geographic locations, consisting of 1.2-second expert driving episodes with 12 frames sampled at , amounting to 1.3 million episodes. The training process was the same as that used in the CARLA experiments. The system trained for 15 days on 64 NVIDIA A100 GPUs. We then deployed the system in closed-loop NVIDIA DRIVE Sim, evaluating its performance across various scenarios and testing its recovery ability by artificially introducing a control lag of approximately . The driving policy demonstrated robust performance and was able to recover effectively. Qualitative results can be seen in supplemental video material.
@@ V CONCLUSION
In this work, we have shown that a driving policy trained with a latent generative world model mitigates the imitation learning covariate shift problem. We also introduced a novel multi-view transformer-based perception encoder using self-supervised pre-training. We showed qualitative and quantitative results on both CARLA and NVIDIA DRIVE Sim. Future work includes enhancing the world model predictor by adopting novel architectures based on linear RNNs or Mamba [37], investigating whether training with the world model reduces the amount of data required for neural policies, and exploring the use of reinforcement learning to handle long-tail scenarios.
@@ Acknowledgment
We thank Arthur Przech, Xiaolin Lin, George Hines, Ephrem Chemali, Ankit Gupta, Seth Robert Piezas, Kayley Ting, Hai Loc Lu, Ryan Holben and Yashraj Narang for contributions, discussions and feedback.

### v3

@@ IV-D DRIVE Sim Experiments
We also tested our model in NVIDIA DRIVE Sim, which is more photo-realistic and physically realistic than CARLA (Fig. 5). We trained our system using a 450-hour internal real driving dataset collected by multiple vehicles in our autonomous vehicle fleet in several geographic locations, consisting of 1.2-second expert driving episodes with 12 frames sampled at , amounting to 1.3 million episodes. The training process was the same as that used in the CARLA experiments. The system trained for 15 days on 64 NVIDIA A100 GPUs. We then deployed the system in closed-loop NVIDIA DRIVE Sim, evaluating its performance across various scenarios and testing its recovery ability by artificially introducing a control lag of approximately . The driving policy demonstrated robust performance and was able to recover effectively. Qualitative results can be seen in supplemental video material. Our model runs in near real time (10Hz) on NVIDIA Orin AGX embedded GPU.
@@ V CONCLUSION
In this work, we have shown that a driving policy trained with a latent generative world model mitigates the imitation learning covariate shift problem. We introduced a novel multi-view transformer-based perception encoder using self-supervised pre-training and DINOv2. We showed qualitative and quantitative results on CARLA Sim, and qualitative results on NVIDIA DRIVE Sim. Future work includes enhancing the world model predictor by adopting architectures based selective state space models like Mamba [38], investigating whether training with the world model reduces the amount of data required for neural policies, and exploring the use of reinforcement learning to handle long-tail scenarios.
@@ Acknowledgment
We thank Arthur Przech, Xiaolin Lin, George Hines, Ephrem Chemali, Ankit Gupta, Seth Robert Piezas, Kayley Ting, Hai Loc Lu, Ryan Holben and Yashraj Narang for contributions and feedback.